# tidychangepoint: a unified framework for analyzing changepoint detection in univariate time series

### v1

@@ 8.1 Future work
One obvious area for improvement in tidychangepoint is computational cost. While the implementations of changepoint detection algorithms in changepoint use pre-compiled C code, all of the code for tidychangepoint is presently written in R. Particularly with respect to genetic algorithms and Bayesian computations, it is likely that pushing some code from R to C could improve performance considerably. Additionally, some algorithms have a known running time in relation to the length of the time series, while others do not.
A second major issue with changepoint detection algorithms is the robustness of the results to a variety of changes in conditions, assumptions, or computational environments. While some algorithms report the probability of an observation being a changepoint, others do not. Some models rely on assumptions about the distribution or frequency of changepoints relative to the length of the time series. In randomized algorithms (such as genetic algorithms), the robustness of the results may be interrelated with stopping criteria, the initial seed or population, or other computational considerations. CUSUM methods operate on a summary of the original time series: are they less susceptible to outliers?

### v2

@@ 8.1 Future work
Generally, computational costs for changepoint detection algorithms are a concern. Some algorithms have a known theoretical running time in relation to the length of the time series, while others do not. Moreover, while the implementations of some algorithms (notably, changepoint and GA) use pre-compiled C code, others may not. All of the code for tidychangepoint is written in R, but since it is mainly an interface that wraps code from other packages, the overhead is slight (see Section 7.2), and there isn’t much that can be done to improve performance. Even the new algorithms that are provided by tidychangepoint (e.g., Coen’s algorithm, Shi’s algorithm) are simply special cases of genetic algorithms that use GA as a backend.
Another major issue with changepoint detection algorithms is the robustness of the results to a variety of changes in conditions, assumptions, or computational environments. While some algorithms report the probability of an observation being a changepoint, others do not. Some models rely on assumptions about the distribution or frequency of changepoints relative to the length of the time series. In randomized algorithms (such as genetic algorithms), the robustness of the results may be interrelated with stopping criteria, the initial seed or population, or other computational considerations. CUSUM methods operate on a summary of the original time series: are they less susceptible to outliers? These issues are not correctable within tidychangepoint. We will continue to engage with the authors of the underlying packages when we see areas for improvement and alert our own users to inconsistencies through our documentation.